they are 9 students buying an antena. they are spliding the amount in equall parts. If 3 more students participate in spliding the amount they would all pay $75 less. How much would the antena cost if they all pay equall amounts?
Answer by jojo14344(1513)

Take a look on this solution:
Let "x" = payment amount of each student
First theory,
There are 9 students so (9)(x) = cost of antenna
Second theory,
If 3 more students joined in, their payment will be $75 less so,
(9+3) (x-75) = cost of antenna
Now, equating the 2 theories since both equals to the cost of the antenna:
9x = (9+3)(x-75)
9x = 12x - 900; 3x = 900; therefore, x =$300 (original equal payment)
When 3 joined in, their equal payment will be x-75= 300-75 = $225/each
Therefore, the cost of the antenna is:
first theory, 9 * 300 = $2700
or
second theory, 12(300-75) = $2700
Very good!